INAF queue observing, 2013 January 28-29

Observer: E. Sani, F. Cusano, r. Carini, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: O Kuhn
Instrument: LBC

Summary

At the beginning of the twilight we have mostly clear sky (thin cirrus just above the NO horizon), with no wind and low humidity (15%).
During the first part of night the seeing is good, ranging between 0.8 and 1.0 on the DIMM. Also the humidity is very low (6%) with no wind
The blu-shutter remained closed during the OB execution. The SA fix the problem turning off and on the camera power from the engineering interface.
At 5.0UT we get problems with the blue channel collimation. We clean the active optics and commit manual corrections for coma.
At 9.0UT thin cirrus passing above N-O horizon.
The shutter problem happened more than once during the night.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Battaglia_DDT IC10blob g-Sloan, i-Sloan 15min, 15 min COMPLETED
Battaglia_DDT IC10fore g-Sloan, i-Sloan 15min, 15 min COMPLETED
Battaglia_DDT IC10ext g-Sloan, i-Sloan 15min, 15 min COMPLETED
Standard SA95 U-Special, B-, V-Bessel, g-, r-Sloan; V-, R-Bessel, r-, i-, z-Sloan - airmass 1.2
Bellazzini_6K HVC196.09 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6A HVC205.28 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Standard SA95 U-Special, B-, V-Bessel, g-, r-Sloan; V-, R-Bessel, r-, i-, z-Sloan - airmass 1.35
Bellazzini_6B HVC204.88 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Giallongo_44A MACSJ0717 U-Special, r-Sloan 30min, 30min within the specs
Bellazzini_6J Leo P g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6Y HVC255.76 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6U HVC250.16 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6L HVC245.26 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6C HVC277.25 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6D HVC274.68 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6M HVC298.95 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6N HVC326.91 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec Executed, but out of seeing specs
Bellazzini_6O HVC28.09 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec Executed. Seeing to be cheked
Bellazzini_6P HVC353.41 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6E HVC351.17 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6Q HVC352.45 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6F HVC356.81 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6R HVC5.58 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED
Bellazzini_6G HVC13.59 g-Sloan, r-Sloan 10min+40sec, 10 min+40sec COMPLETED

Details

00:46 The TO opens the dome.

01:17 Start with flats calibrations
  • Run SkyFlat _SV_r0429d+54_pa0_5.ob R=012037 has low counts
  • Run SkyFlat _SV_r0429d+54_pa180_5.ob with scaling factor 4.0. Good in STD, low counts in r

01:36 Slew to Battaglia_DDT IC10
  • Run IC10blob_copoint.ob
  • dofpia, /first, /X2
  • lbcrangebal B=14640 R=14632

01:49 Run IC10blob_gisloan_1.ob
  • the average seeing at the beginning of the integration is 1.0 in the red channel and 1.2 in the blue one.

02:10 Run IC10fore_gisloan_2.ob
  • dofpia, /X2
  • we do not perform co-pointing because the fields are very close. * the average seeing at the beginning of the integration is 1.0 in both channel.

02:35 Run IC10ext_gisloan_3.ob
  • dofpia, /X2
  • we do not perform co-pointing because the fields are very close.
  • the average seeing at the beginning of the integration is 1.0 in both channel.

03:00 Slew to the standard Slew to the standard SA95
  • dofpia, /X2

03:27 Slew to Bellazzini_6K HVC196.09+24.74+166
  • Run 6_K_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=033332 R=033329

03:38 Run 6_K.ob
  • the average seeing is 1.0 in both channels.

03:53 Slew to Bellazzini_6A HVC205.28+18.70+150
  • Run 6_A_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=040127 R=040123

04:04 Run 6_A.ob
  • the average seeing is 0.8 in both channels.
  • Little elongation seen in the short exposure images of the blue channel. The profile of stars ahs anyhow low dispersion, we thus consider this OB done.

04:21 Slew to Bellazzini_6B HVC204.88+44.86+147
  • the target is still too low.

04:24 Slew to the standard Slew to the standard SA95
  • dofpia, /X2
  • the seeing is still below 1arcsec, we see some elongation in blue images

04:44 Problems with blue images: there is no signal in B images 044125, 044205, 044245, 044326, 044413, 044454
  • the shutter remained closed during the OB execution. The SA fix the problem turning off and on the camera power from the engineering interface.
  • after a test-image check, we repeat the SA95 integration

04.59 Run SA95_SBVgr_VRriz.ob

05:10 Slew to Bellazzini_6B HVC204.88+44.86+147
  • Run 6_B_copoint.ob
  • dofpia, /X2. There are very few pupils especially in the blue channel. We slew to the scientific target to collimate.
  • There is a strong coma on the blue channel, the TO operates corrections by hand. It doesn't work. We ask the TO to clear the adaptive optics.

05:44 We commit manual corrections +2000 for Z7 and -2000 for Z8 fo the blu channel.
  • Run dofpia, /first, /X2, /bleonly on the scientific field Bellazzini_6B
  • PA wrong, commit -3000 for Z7
  • Run dofpia, /first, /X2, /bleonly on the scientific field Bellazzini_6B. It doesn't work. Stop the script and commit again manual corrections: +1000 for both Z7 and Z8
  • Run dofpia, /first, /X2, /bleonly

06:03 Run 6_B.ob
  • the average seeing is 1.0 in the blue channel and 0.8 in the red one.

06:20 Slew to Giallongo_44A MACSJ0717
  • Run 44A_g_r_pointing_1.ob
  • dofpia, /X2
  • lbcrangebal B=63021 R=63015

06:36 Run 44A_1.ob
  • the average seeing is 1.0 in the blue channel and 0.9 in the red one.

07:16 Slew to Bellazzini_6J Leo P
  • Run 6_J_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=72543 R=72537

07:30 Run 6_J.ob
  • the average seeing is 0.7 in the blue channel and more 0.6 in the red one.

07:48 Slew to Bellazzini_6Y HVC255.76+61.49+181
  • Run 6_Y_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=75040 R=75034

07:54 Run 6_Y.ob
  • the average seeing is 0.6 blue 0.7 red channels.

08:09 Slew to Bellazzini_6U HVC250.16+57.45+139
  • Run 6_U_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=81810 R=81805

08:21 Run 6_U.ob
  • the average seeing is 1.1 blue 0.8 red channels.

08:49 Slew to Bellazzini_6L HVC245.26+69.53+217
  • since the collimation is stable, we skip dofpia. The target has the same elevation than 6U, thus we also skip the copointing
  • Run 6_L.ob
  • the average seeing is 1.1 blue 0.8 red channels.

08:51 Slew to Bellazzini_6C HVC277.25+65.14-140
  • Run 6_C_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=85730 R=85726

09:00 Run 6_C.ob
  • the average seeing is 0.9 on both channels

09:15 Slew to Bellazzini_6D HVC274.68+74.70-123
  • since the collimation is stable, we skip dofpia. The target has the same elevation than 6C, thus we also skip the copointing
  • Run 6_D.ob
  • image lbcb.092928 has no signal, we get the shutter stucked again. The SA switch-off and -on the camera. We take a test image, it is fine

09:36 We run again 6_D.ob integrating only the 4s images (in both channels)
  • the seeing is 1.2 in blue and 1.0 in red

09:41 Slew to Bellazzini_6M HVC298.95+68.17+270
  • Run 6_M_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=94705 R=94701

09:48 Run 6_M.ob * the average seeing is 1.0 on both channels.

10:05 Slew to Bellazzini_6N HVC326.91+65.25+316
  • since the collimation is stable, we skip dofpia. The target has the same elevation than 6M, thus we also skip the copointing
  • Run 6_N.ob
  • the seeing is 1.4 in blue and 1.2 in red. Out of seeing specs (to be confirmed by allseeing)

10:19 Slew to Bellazzini_6O HVC28.09+71.86-144
  • Run 6_O_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=102720 R=102718

10:30 Run 6_O.ob * We have to re-run the scientific OB because there is a filter error. OB executed, there was some glitch between the telescope and the instrument * Possibly the seeing in the blue channel is out of specs

10:49 Slew to Bellazzini_6P HVC353.41+61.07+257 * Run 6_P_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=105523 R=105518

11:00 Run 6_P.ob
  • the seeing is 0.8 in both channels

11:15 Slew to Bellazzini_6E HVC351.17+58.56+214
  • since the collimation is stable, we skip dofpia. The target has the same elevation than 6p, thus we also skip the copointing
  • the seeing is 1.2 in blue and 1.0 in red

11:31 Slew to Bellazzini_6Q HVC352.45+59.06+263
  • Run 6_Q_copoint.ob
  • dofpia, /X2
  • lbcrangebal B=113729 R=113725

11:42 Run 6_Q.ob
  • At the beginning of the integration there are cirrus passing through. This should affect only the first 5min frames.
  • the average seeing is 0.9 in blue and 1.0 in red
  • some elongation in the red image lbcr.114803. Note: dofpia converged with only two pupils in the red channel.

11:57 Slew to Bellazzini_6F HVC356.81+58.51+148
  • run dofpia on scientific field Bellazzini_6F

12:02 Run 6_F.ob
  • the seeing is 0.9in both the channel red and blue

12:17 Slew to Bellazzini_6R HVC5.58+52.07+163
  • dofpia, /X2 on scientific fiels 6_R

12:24 Run 6_R.ob

12:50 Slew to Bellazzini_6G HVC 13.59+54.52+169
  • Run 6_G_point.ob
  • dofpia, /X2
  • lbacrangebal

12:55 Run 6_G the seeing is 0.8 in blue and 0.7 in red

13:05 Slew to Bellazzini_6H HVC 13.60+54.23+179
  • since the collimation is stable, we skip dofpia. The target has the same elevation than 6G, thus we also skip the copointing
  • we are actually out of focus. we thus run dofpia on the scientific field Bellazzini_6H
  • backgound count are twice the prvious images. The twilight is already too bright

13:18 We start flats calibration
  • Run SkyFlat _gi_r1502d+29_pa0_5.ob with scaling 2 and than 1
  • Run SkyFlat _gi_r1502d+29_pa180_5.ob with scaling 1 and 0.3. The shutter was stucked for the image lbcr.134913

13.51 The TO closes the dome
  • Some test to solve the shutter issue.

14:213 Run 25BIAS_BINO.ob

-- EleanoraSani - 29 Jan 2014
Topic revision: r5 - 20 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback